### v2

@@ Abstract
Medical image segmentation is a pivotal step in diagnostic and therapeutic processes, relying on high-quality annotated data that is often challenging and costly to obtain. Semi-supervised learning offers a promising approach to enhance model performance by leveraging unlabeled data. Although weak-to-strong consistency is a prevalent method in semi-supervised image segmentation, there is a scarcity of research on perturbation strategies specifically tailored for semi-supervised medical image segmentation tasks. To address this challenge, this paper introduces a simple yet efficient semi-supervised learning framework named Attention-Guided weak-to-strong Consistency Match (AIGCMatch). The AIGCMatch framework incorporates attention-guided perturbation strategies at both the image and feature levels to achieve weak-to-strong consistency regularization. This method not only preserves the structural information of medical images but also enhances the model’s ability to process complex semantic information. Extensive experiments conducted on the ACDC and ISIC-2017 datasets have validated the effectiveness of AIGCMatch. Our method achieved a 90.4% Dice score in the 7-case scenario on the ACDC dataset, surpassing the state-of-the-art methods and demonstrating its potential and efficacy in clinical settings. Additionally, on the ISIC-2017 dataset, we significantly outperformed our baseline, indicating the robustness and generalizability of AIGCMatch across different medical image segmentation tasks.
Keywords:
@@
Weak-to-Strong Consistency Medical Image Segmentation Semi-Supervised Learning.
1 Introduction
@@
Medical image segmentation is a fundamental task aimed at delineating anatomical structures such as organs and tumors by classifying each pixel into distinct categories [32, 17]. The precision achieved through medical image segmentation provides invaluable volumetric and shape details, serving as foundational elements for disease diagnosis, treatment efficacy evaluation, and quantitative analysis within clinical settings[9, 24, 14]. Data-driven approaches, particularly those based on fully supervised learning, have shown remarkable performance in this domain [26, 8, 22]. However, their success relies heavily on the availability of large-scale, high-quality annotated datasets, which are often challenging and expensive to obtain [31, 17]. This limitation has spurred the development of semi-supervised learning (SSL) approaches, which leverage abundant unannotated data alongside a smaller set of annotated examples [23, 39].
In the domain of SSL, consistency regularization has emerged as a key technique to use unannotated data to improve model performance  [23, 5, 16, 20, 33, 45]. The principle behind consistency regularization requires the model to produce consistent predictions on different perturbations of the input data. This is particularly relevant in medical imaging, where slight variations in image acquisition parameters or patient positioning can lead to variations in the appearance of the same anatomical structures  [17]. Despite the remarkable performance of consistency Regularization, a critical limitation in current methods is their reliance on random perturbations to simulate data variability  [37]. This approach, while effective for natural images with clear foreground-background distinctions, may not fully capture the complex high-dimensional semantic information of medical images. For example, in Fig 1, a model can easily segment a cat from the background in a scene because of the significant variation in pixels. However, segmenting organs or tissues from images not only requires the injection of proprietary semantic information but there are no clear pixel differences between different categories. This discrepancy between the natural images and the medical images calls for a specialized perturbation strategy tailored for semi-supervised medical image segmentation, a gap that existing works have yet to adequately address.
This realization has catalyzed the development of our novel framework, AttentIon -Guided Consistency regularization Match (AIGCMatch), which rethinks the perturbation process. The framework shifts the focus from random perturbations to attention-guided perturbations. By harnessing the power of attention mechanisms, our method rethinks the notions of strong and weak consistency regularization at both the image and feature levels. At the feature level, our method leverages the skip connection structure of the U-net, utilizing the channel attention mechanism generated by the encoder layers to identify important channels within the feature maps. We introduce noise into these salient channels, as well as into less significant ones, allowing the features with perturbation to be passed along to the corresponding decoder layers through the U-net’s skip connections. This process enables the model to supervise the predictions from the noised important channels input using the predictions from the noised unimportant channels input. The method of consistency regularization at the feature level is called AttFeaPerb in our paper. At the image level, we depart from previous consistency regularization methods that employed random perturbation techniques such as CutMix [38]. Instead, we utilize the model’s attention map to identify key regions within the image for targeted CutMix as a strong perturbation named attention-guided CutMix (AttCutMix). We then use weak perturbations at the image level to supervise the predictions from the input. In summary, unlike random perturbations, attention-guided perturbations have two advantages: 1) At the image level, attention-guided CutMix perturbs semantically related areas as a whole, maintaining the perturbation intensity without destroying the structural information of medical images. 2) At the feature level, perturbing by selecting channels with high attention is conducive to generating stronger perturbations to the model, promoting more robust model predictions. Our contributions are summarised as follows:
@@
- Through extensive experimentation on the ACDA dataset, we validate the efficacy of AIGCMatch. Our framework consistently outperforms state-of-the-art methods, demonstrating its effectiveness and potential for widespread adoption in clinical settings.
2 Related Work
2.1 Medical Image Analysis
@@
Medical image analysis is crucial for computer-aided diagnosis and treatment planning, with applications in identifying pathological tissues and tracking disease progression [12]. Advances in imaging modalities like CT, MRI, and US have significantly enhanced the visualization of internal structures. A pivotal task in this field is medical image segmentation, which aids in precise diagnoses and planning by delineating regions of interest. The advent of deep learning, particularly Encoder-Decoder architectures, has revolutionized medical image segmentation. Early network design breakthroughs included the U-Net [26], which introduced a symmetric encoder-decoder structure and skip connections, setting new standards for accuracy and efficiency. Building on U-Net’s success, variants like U-Net++ [44] and U-Net3+ [10] were proposed to improve feature fusion and segmentation performance with additional enhancements. The emergence of the Visual Transformer (ViT) model has led to hybrid architectures combining CNNs for local spatial information and Transformers for long-range dependencies and semantic processing [41, 34, 21].
2.2 Semi-supervised medical image segmentation
@@
However, medical image segmentation based fully supervised are heavily dependent on large volumes of expert-annotated data, which is a time-consuming, labor-intensive, and expensive process [15]. Additionally, these models may suffer from reduced effectiveness when encountering novel data or shifts in data distribution, thereby constraining their generalizability[16].To address the above-mentioned question, researchers have introduced semi-supervised learning into medical image segmentation.Previous methods of semi-supervised medical image segmentation rely on the combination of a limited amount of labeled data and a vast quantity of unlabeled data to enhance model performance. These methods primarily include pseudo-labeling, graphical models, generative models, and autoencoders.
Pseudo-labeling involves utilizing predictions from a model trained on labeled data to generate pseudo-labels for unlabeled data. These pseudo-labels are then used to augment the training set and train the model iteratively[40, 35, 19].Graphical model approaches to leverage the structure of graphs to represent pixel relationships within images, maximizing the posterior probability of labeled data for segmentation[29, 1, 6]. As highlighted by Litjens et al.[17], these methods typically construct an energy function that models the relationships between pixel labels and neighboring pixels, promoting label consistency within similar regions. Although these approaches perform well on small-scale datasets, they often face computational complexity and a lack of labeled information when applied to large-scale imaging data.Generative models, particularly Generative Adversarial Networks (GANs), have also gained widespread application in semi-supervised medical image segmentation[30, 13, 11]. Ouali et al.[25] note that GANs, through the adversarial training of generators and discriminators, not only generate realistic samples but also effectively leverage unlabeled data. This capability enables models to achieve improved segmentation performance with limited labeled samples. Autoencoder is another commonly used method in semi-supervised learning[7, 27, 28], which compresses input images into lower-dimensional representations and then reconstructs the original images. As mentioned by Sajjadi et al. [27], through self-supervised learning on unlabeled data, autoencoders can extract important features that provide richer information for segmentation tasks.
@@
While these previous methods laid the foundation for subsequent techniques, they still encounter challenges in dealing with imbalanced data and noise in complex scenarios.
2.3 Applications of Consistency Regularization and Weak-to-Strong Consistency in Semi-Supervised Medical Image segementation
Consistency regularization[5, 3, 20] serves as an effective semi-supervised learning strategy that enhances the robustness of models to input perturbations, thus advancing medical image segmentation. According to Xie et al.[36], this strategy ensures that the model’s predictions remain consistent across various transformations of the input data, thereby improving generalization and providing new perspectives for the effective utilization of unlabeled data.
In practical applications, strong consistency and weak consistency are two essential components of consistency regularization, each playing a distinct role. Strong consistency mandates that the model maintains output consistency under significant transformations, such as rotation, scaling, or color changes. This strategy aids in enhancing the model’s stability in complex scenarios, ensuring that it can still make accurate predictions in the face of real-world variations. Research by Ouali et al.[25] indicates that imposing strong consistency constraints can significantly improve model performance on complex medical images.
In contrast, weak consistency focuses on maintaining prediction consistency under minor perturbations, such as low noise levels or slight geometric changes. Sajjadi et al.[27] emphasize that this strategy effectively captures subtle structural variations, making it particularly suitable for complex biomedical images. Weak consistency not only enhances the model’s resistance to noise but also improves segmentation accuracy, rendering semi-supervised learning more practical in medical image segmentation tasks.
In previous works, the effectiveness of strong consistency and weak consistency in image segmentation has been well established. However, when it comes to weak-to-strong consistency, applications of this simple yet effective method in the field of medical imaging remain relatively scarce. Yang et al. proposed Unimatch [37], a novel framework for semi-supervised image segmentation based on weak-to-strong consistency, which achieved state-of-the-art performance across various datasets, including the ACDC dataset. This achievement highlights the potential of strong and weak consistency in semi-supervised medical image segmentation and has inspired a series of subsequent works, such as CrossMatch[42]. Inspired by previous work, we propose an attention-guided perturbation approach that emphasizes contextual information in medical images and improves the ability of models for semi-supervised medical image segmentation.
@@
3 Method
3.1 The Overview of Framework
As depicted in Fig2, the proposed framework consists of two attention-guided perturbation modules in image and feature level, aiming to enhance the capability of the semi-supervised segmentation model to capture complex structural and high-dimensional semantic information in medical images. To enhance the pertinence of perturbations to the model, we fully leverage the attention from various parts of the model. At each level, the perturbations are crafted according to the attention derived from the respective parts of the model. In Fig 3, we compared the differences between our model and the Unimatch[37]. It comprises a supervised branch and two unsupervised branches. The model is supervised for labeled data by computing the loss between the model’s predictions and the ground truth (GT). For unlabeled data, attention maps from the model are employed to guide perturbations at both the feature and the image levels. Specifically, at the feature level, we proposed AttFeaPerb to obtain strong perturbation predictions $p^{fsp}\in\mathbb{R}^{C\times H\times W}$ and weak perturbation predictions $p^{fwp}\in\mathbb{R}^{C\times H\times W}$. At the image level, we used predictions $p^{w}\in\mathbb{R}^{C\times H\times W}$ obtained from weak perturbations to supervise the predictions $p^{s1}\in\mathbb{R}^{C\times H\times W}$ and $p^{s2}\in\mathbb{R}^{C\times H\times W}$ obtained from AttCutMix.
3.2 Attention-Guided Perturbation in Image Level
@@
3.3 Exploring Weak-to-Strong Consistency in the Feature Level
@@ 3.4 Loss Function
$\mathcal{L}_{u}=\frac{1}{B_{u}}\sum\mathbbm{1}\left(\max(p^{w})\geq\tau\right)\cdot\left(\frac{\lambda}{2}\left(\mathrm{H}(p^{w},p^{s_{1}})\right)+\mathrm{H}(p^{w},p^{s_{2}})\right)$ (11)
$+\mathbbm{1}\left(\max(p^{fwp})\geq\tau\right)\cdot\left(\mu\mathrm{H}(p^{fwp},p^{fsp})\right)$
@@
The weights for the loss function, denoted as $\lambda$ and $\mu$, are set to 0.5. In the unsupervised loss $L_{u}$, the function $\mathrm{H}$ is commonly used as the Dice coefficient loss.
4 Experiments
@@
Figure 4: Comparison with ground truth and two different results. Three different colors depicted in the figure represent three distinct segmentation outcomes. The section enclosed in a light blue box indicates areas where the model has either made an erroneous prediction or has omitted a prediction.
4.1 Experimental Setups
4.1.1 Datasets
@@
In this research, we evaluated our model’s performance on the ACDC and ISIC-2017 within a semi-supervised learning framework for medical image segmentation. The ACDC dataset[4], consisting of 100 two-dimensional cine cardiac MR images from the University Hospital of Dijon, is divided into 70 scans for training, 10 for validation, and 20 for testing. It is unique in targeting multiclass segmentation of cardiac structures, contrasting with other datasets that focus on binary segmentation. Within the domain of semi-supervised medical image segmentation, the ACDC dataset is conventionally divided into distinct segments categorized as 1-case, 3-case, and 7-case. These categories correspond to the utilization of 1%, 5%, and 10% of the available labeled data, respectively.
For the ISIC-2017 dataset, we adopted a similar semi-supervised learning strategy. This dataset, which encompasses a wide range of skin lesion images, is an excellent resource for evaluating segmentation models in the context of dermatology. We utilized 10% of the labeled images for training, which amounts to 200 images, and the remaining 90% as unlabeled data for our semi-supervised model. The validation set within ISIC-2017 contains 150 images, which were used to evaluate the model’s performance in a real-world scenario.
4.1.2 Implementation Details

### v3

@@ Abstract
Medical image segmentation is a pivotal step in diagnostic and therapeutic processes, relying on high-quality annotated data that is often challenging and costly to obtain. Semi-supervised learning offers a promising approach to enhance model performance by leveraging unlabeled data. Although weak-to-strong consistency is a prevalent method in semi-supervised image segmentation, there is a scarcity of research on perturbation strategies specifically tailored for semi-supervised medical image segmentation tasks. To address this challenge, this paper introduces a simple yet efficient semi-supervised learning framework named Attention-Guided weak-to-strong Consistency Match (AIGCMatch). The AIGCMatch framework incorporates attention-guided perturbation strategies at both the image and feature levels to achieve weak-to-strong consistency regularization. This method not only preserves the structural information of medical images but also enhances the model’s ability to process complex semantic information. Extensive experiments conducted on the ACDC dataset have validated the effectiveness of AIGCMatch. Our method achieved a 90.4% Dice score in the 7-case scenario on the ACDC dataset, surpassing the state-of-the-art methods and demonstrating its potential and efficacy in clinical settings.
Index Terms:
@@
Weak-to-Strong Consistency, Medical Image Segmentation, Semi-Supervised Learning.
footnotetext: Equal contribution. Corresponding author.
I Introduction
@@
Medical image segmentation is a fundamental task aimed at delineating anatomical structures such as organs and tumors by classifying each pixel into distinct categories [1, 2]. The precision achieved through medical image segmentation provides invaluable volumetric and shape details, serving as foundational elements for disease diagnosis, treatment efficacy evaluation, and quantitative analysis within clinical settings [3, 4]. Data-driven approaches, particularly those based on fully supervised learning, have shown remarkable performance in this domain [5, 6, 7]. However, their success relies heavily on the availability of large-scale, high-quality annotated datasets, which are often challenging and expensive to obtain [2]. This limitation has spurred the development of semi-supervised learning (SSL) approaches, which leverage abundant unannotated data alongside a smaller set of annotated examples [8, 9].
In the domain of SSL, consistency regularization has emerged as a key technique to use unannotated data to improve model performance [8, 10, 11, 12, 13, 14]. The principle behind consistency regularization requires the model to produce consistent predictions on different perturbations of the input data. This is particularly relevant in medical imaging, where slight variations in image acquisition parameters or patient positioning can lead to variations in the appearance of the same anatomical structures [2]. However, despite the remarkable performance of consistency regularization in general computer vision tasks, its direct application to medical image segmentation faces two significant challenges that limit its effectiveness in clinical applications:
C1. How can we design perturbation strategies that respect the unique characteristics of medical images? Existing approaches predominantly rely on random perturbation techniques originally designed for natural images, which fail to account for the complex anatomical structures present in medical data [15]. As illustrated in Fig. 1, natural images typically exhibit distinct pixel-level differences between objects and backgrounds, making them amenable to random perturbation techniques. For instance, a cat can be easily differentiated from its background due to clear visual boundaries and color contrasts. In contrast, medical images contain complex anatomical structures with subtle intensity variations between different tissues, where adjacent regions may appear visually similar yet represent entirely different anatomical entities. These subtle differences require preservation of specific structural and contextual information that random perturbations may inadvertently disrupt. Consequently, generic perturbation strategies may compromise the semantic integrity essential for accurate medical image segmentation, highlighting the need for perturbation approaches specifically tailored to the unique characteristics of medical imaging data.
C2. How can we extend the scope of perturbations beyond the image level to capture high-dimensional semantic information? Current methods predominantly focus on perturbations at the image level, overlooking the potential of feature-level perturbations to enhance model robustness. Medical images contain complex high-dimensional semantic information that is progressively extracted and refined through different layers of deep learning models. By restricting perturbations to the input level, these methods fail to challenge the model’s ability to maintain consistent representations across different feature abstraction levels. This limitation is particularly significant in medical image segmentation, where subtle feature variations can substantially impact segmentation boundaries. A comprehensive approach that incorporates perturbations at multiple levels of representation could potentially improve the model’s capacity to capture and maintain consistent semantic information throughout the segmentation process.
To address these challenges, we propose AttentIon-Guided Consistency regularization Match (AIGCMatch), a novel semi-supervised framework for medical image segmentation. ❶ To address the first challenge of inappropriate perturbation strategies, we introduce attention-guided CutMix (AttCutMix), which leverages the model’s attention mechanisms to identify semantically meaningful regions for targeted perturbations. Unlike random perturbations such as CutMix [16], AttCutMix preserves the structural integrity of medical images while introducing beneficial variability, ensuring that perturbations respect anatomical boundaries and tissue relationships. ❷ To address the second challenge of limited perturbation scope, we propose attention-guided feature perturbation (AttFeaPerb), a novel technique that extends consistency regularization to the feature level by selectively perturbing channels with varying attention scores. This approach establishes a weak-to-strong consistency paradigm at the feature level, enhancing the model’s ability to maintain robust representations across different levels of abstraction. By integrating these attention-guided perturbation strategies at both image and feature levels, AIGCMatch effectively captures the complex structural and high-dimensional semantic information in medical images, significantly improving segmentation performance while reducing the requirement for labeled data.
Our contributions are summarized as follows:
@@
  • Through extensive experimentation on the ACDC dataset, we validate the efficacy of AIGCMatch. Our framework consistently outperforms state-of-the-art methods, demonstrating its effectiveness and potential for widespread adoption in clinical settings.
II Related Work
II-A Medical Image Analysis
@@
Medical image analysis is crucial for computer-aided diagnosis and treatment planning, with applications in identifying pathological tissues and tracking disease progression [17]. Advances in imaging modalities like CT, MRI, and US have significantly enhanced the visualization of internal structures. A pivotal task in this field is medical image segmentation, which aids in precise diagnoses and planning by delineating regions of interest. The advent of deep learning, particularly Encoder-Decoder architectures, has revolutionized medical image segmentation. Early network design breakthroughs included the U-Net [5], which introduced a symmetric encoder-decoder structure and skip connections, setting new standards for accuracy and efficiency. Building on U-Net’s success, variants like U-Net++ [7] and U-Net3+ [18] were proposed to improve feature fusion and segmentation performance with additional enhancements. The emergence of the Visual Transformer (ViT) model has led to hybrid architectures combining CNNs for local spatial information and Transformers for long-range dependencies and semantic processing [19, 20].
II-B Semi-supervised medical image segmentation
@@
However, medical image segmentation based fully supervised are heavily dependent on large volumes of expert-annotated data, which is a time-consuming, labor-intensive, and expensive process [21]. To address the question, researchers have introduced semi-supervised learning into medical image segmentation, and they primarily include pseudo-labeling, generative models, and autoencoders.
Pseudo-labeling involves utilizing predictions from a model trained on labeled data to generate pseudo-labels for unlabeled data. These pseudo-labels are then used to augment the training set and train the model iteratively [22, 23, 24].
Generative models, particularly Generative Adversarial Networks (GANs), have also gained widespread application in semi-supervised medical image segmentation[25, 26, 27]. Ouali et al.[28] note that GANs, through the adversarial training of generators and discriminators, not only generate realistic samples but also effectively leverage unlabeled data. This capability enables models to achieve improved segmentation performance with limited labeled samples.
Autoencoder is another commonly used method in semi-supervised learning[29, 30], which compresses the input images into lower-dimensional representations and then reconstructs the original images. As mentioned by Sajjadi et al. [30], through self-supervised learning on unlabeled data, autoencoders can extract important features that provide richer information for segmentation tasks.
@@
Consistency Regularization While these previous methods laid the foundation for subsequent techniques, they still encounter challenges in dealing with imbalanced data and noise in complex scenarios.
Consistency Regularization [10, 31, 12] is the most popular technique in semi-supervised medical image segmentation. Many works focus on ensuring the model outputs the same result under different perturbations. Basak et al. propose a method of consistency prediction using interpolated unlabeled images to improve the performance of semi-supervised medical image segmentation[31]. Luo et al. proposed a semi-supervised medical image segmentation framework leveraging dual-task consistency to improve segmentation accuracy by jointly predicting pixel-wise segmentation and geometry-aware level representations[12].
@@
III Method
III-A The Overview of Framework
As depicted in Fig 2, the proposed framework consists of two attention-guided perturbation modules in the image and feature level, aiming to enhance the capability of the semi-supervised segmentation model to capture complex structural and high-dimensional semantic information in medical images. To enhance the pertinence of perturbations to the model, we fully leverage the attention from various parts of the model. At each level, the perturbations are crafted according to the attention derived from the respective parts of the model. In Fig 3, we compared the differences between our model and the Unimatch[15]. It comprises a supervised branch and two unsupervised branches. The model is supervised for labeled data by computing the loss between the model’s predictions and the ground truth (GT). For unlabeled data, attention maps from the model are employed to guide perturbations at both the feature and the image levels. Specifically, at the feature level, we proposed AttFeaPerb to obtain strong perturbation predictions $p^{fsp}\in\mathbb{R}^{C\times H\times W}$ and weak perturbation predictions $p^{fwp}\in\mathbb{R}^{C\times H\times W}$. At the image level, we used predictions $p^{w}\in\mathbb{R}^{C\times H\times W}$ obtained from weak perturbations to supervise the predictions $p^{s1}\in\mathbb{R}^{C\times H\times W}$ and $p^{s2}\in\mathbb{R}^{C\times H\times W}$ obtained from AttCutMix.
III-B Attention-Guided Perturbation in Image Level
@@
III-C Exploring Weak-to-Strong Consistency in the Feature Level
@@ III-D Loss Function
[…]
@@
In the unsupervised loss $L_{u}$, the function $\mathrm{H}$ is commonly used as the Dice coefficient loss.
IV Experiments
@@
Figure 4: Comparison with ground truth and two different results. The three different colors depicted in the figure represent three distinct segmentation outcomes. The section enclosed in a light blue box indicates areas where the model has either made an erroneous prediction or omitted a prediction.
Method Reference 1 case 3 cases 7 cases
[…]
SupBaseline 28.50 16.62 47.83 37.01 31.16 12.62 79.41 68.11 9.35 2.70
UA-MT [33] MICCAI’19 42.28 32.21 40.74 18.58 72.71 60.89 21.48 7.15 85.14 75.90 6.25 1.80
ICT [34] IJCAI’19 43.03 30.58 34.92 15.23 74.20 62.90 17.01 4.32 85.15 76.05 4.27 1.46
SASSNet [35] MICCAI’20 57.77 46.14 20.05 6.06 84.50 74.34 5.42 1.86
FixMatch [36] NeurIPS’20 69.67 58.34 37.92 14.41 83.12 73.59 9.86 2.61 88.31 79.97 7.35 1.79
DTC [12] AAAI’21 56.90 45.67 23.36 7.39 84.29 73.92 12.81 4.01
CPS [37] CVPR’21 56.70 44.31 24.97 10.48 75.24 64.67 10.93 2.98 84.63 75.20 7.57 2.27
MCNetV2 [38] MIA’22 57.49 43.29 31.31 10.97 78.96 68.15 12.13 3.91 85.97 77.21 7.55 2.11
MC-Net [39] MICCAI’21 62.85 52.29 7.62 2.33 86.44 77.04 5.50 1.84
URPC [40] MedIA’22 55.87 44.64 13.60 3.74 83.10 72.41 4.84 1.53
SS-Net [41] MICCAI’22 65.82 55.38 6.67 2.28 86.78 77.67 6.07 1.40
BCP [42] CVPR’23 87.59 78.67 1.90 0.67 88.84 80.62 3.98 1.17
DMD [43] MICCAI’23 80.60 69.08 5.96 1.90 87.52 78.62 4.81 1.60
INCL [44] MIDL’23 77.80 66.13 11.69 3.22 85.43 75.76 6.37 1.37 88.28 80.09 1.67 0.49
Unimatch [15] CVPR’23 85.43 74.57 88.86 79.95 89.85 81.57
URCA [45] CMPB’24 83.31 6.95 2.16 87.86 4.21 1.36
CrossMatch [46] arXiv’24 88.27 80.17 1.53 0.46 89.08 81.44 1.52 0.52
DiffRect [47] MICCAI’24 82.40 71.96 10.04 2.90 86.95 78.08 4.07 1.23 90.18 82.12 1.38 0.48
Ours 86.25±2.50 75.82 1.11 0.21 89.53±2.15 81.07 0.94 0.17 90.40±1.68 82.48 0.86 0.14
TABLE I: Comparison with different models on the ACDC dataset, using 1, 3, and 7 labeled cases. Methods are ordered chronologically by publication year.
IV-A Experimental Setups
IV-A1 Datasets
@@
In this research, we evaluated our model’s performance on the ACDC within a semi-supervised learning framework for medical image segmentation. The ACDC dataset[48], consisting of 100 two-dimensional cine cardiac MR images from the University Hospital of Dijon, is divided into 70 scans for training, 10 for validation, and 20 for testing. It is unique in targeting multiclass segmentation of cardiac structures, contrasting with other datasets that focus on binary segmentation. Within the domain of semi-supervised medical image segmentation, the ACDC dataset is conventionally divided into distinct segments categorized as 1-case, 3-case, and 7-case. These categories correspond to the utilization of 1%, 5%, and 10% of the available labeled data, respectively.
IV-A2 Implementation Details